Can someone help me solve this?
Solve the system by the elimination method.
3x-5y=1
6x-10y=4

Answer by stanbon(75887)

Solve the system by the elimination method.
3x-5y=1
6x-10y=4
------
Modify for elimination::
6x - 10y = 2
6x - 10y = 4
-------------
Subtract to get:
0 = 2
This is a contradiction and implies
there is no (x,y) that satisfies the system.
================
Cheers,
Stan H.
-----------